# Assembly Theory Reduced to Shannon entropy and Rendered Redundant by Naive Statistical Algorithms

### v7

@@ Abstract
We respond to arguments against our critique of Assembly Theory that claim to show its divergence from popular compression. We have proven that any implementation of the concept of ‘copy number’ underlying Assembly Theory (AT) and its assembly index (Ai) is equivalent to Shannon entropy, and not fundamentally or methodologically different from algorithms like ZIP compression. Here we show that the weak empirical correlation between Ai and LZW, which the authors have offered as a defence is based on an incomplete and misleading experiment. When the experiment is completed, the fast asymptotic convergence to $\mathbf{LZ}$ compression and Shannon entropy is undeniable, just as their mathematical proof of equivalence remains undisputed. This contribution completes the theoretical and empirical demonstration that any variation of the copy-number concept underlying AT, which entails counting the number of object repetitions ‘to arrive at a measure for life,’ is equivalent to statistical compression and Shannon entropy. We demonstrate that the authors’ ‘we-are-better-because-we-are-worse’ argument against compression does not withstand basic scrutiny, and that their empirical results separating organic from inorganic compounds have not only been previously reported, sans claims to unify physics and biology, but are driven solely by molecular length– which they did not control for– rather than by any particular feature of life captured by Assembly Theory or their index. Finally, we show that Ai is a particular case of our Block Decomposition Method, introduced almost a decade earlier, and that arguments attributing special stochastic properties to Ai are misleading, since the properties of Ai are not unique, but rather exactly the same as those that Shannon entropy possesses and for which it was designed in the first place–for the quantification of uncertainty–and which we have proven to be equivalent to Ai. Shannon entropy is already not only equipped with stochasticity but was designed for. This makes AT redundant especially when applied to their own experimental data on which we have proven both Shannon entropy and LZW compression deliver the same or even better results that we further expanded here. We will show that contrary to their claims, arguments related to time computational complexity as a proof of dissimilarity are mistaken.
keywords:
@@
Biosignatures, compression, Assembly Theory, Algorithmic Complexity, Shannon entropy, LZ, LZW, computability, stochasticity, invariability, data representation.
1 Introduction
In a recent article responding to some of our criticisms, purported evidence was offered that a complexity measure called the ‘assembly index’, central to the arguments of a hypothesis called Assembly Theory, was not a popular lossless compression, specifically not LZW [1] introduced by Lempel, Ziv and Welch in the late 1970s.
Earlier findings demonstrated that the Assembly Theory (AT) and its assembly index (Ai), or molecular assembly (MA), when applied in the context of molecular complexity, performed similarly—and sub-optimally—when compared to simple statistical compression algorithms like LZW, with a strong Pearson correlation of 0.874 between LZW and MA [2]. It was also shown in the same work that LZW and other compression methods, including those based on Shannon entropy ($\mathbf{H}$), were able to tell apart organic from inorganic compounds using mass spectral data as effectively as, or better than, the assembly index. We concluded that Ai [2], and therefore AT, offered no substantial improvement over existing statistical methods for identifying biosignatures or separating living from non-living systems, which has been a central basis of the Assembly Theory authors’ claim that their theory explained evolution and selection, and unified biology and physics [3].
We also provided a mathematical proof that the assembly index is a lossless compression scheme based on the Lempel-Ziv factorisation family of algorithms, and that such a scheme, to which the assembly index calculation method is equivalent, cannot outperform $\mathbf{H}$ even in the best case scenario [4], hence making it indistinguishable from $\mathbf{LZ}$ compression and Shannon entropy itself.
@@ 1 Introduction
The assembly index (Ai) implements Assembly Theory’s (AT’s) main hypothesis, viz. that the number of exact copies in an object is a measure of its selective content [3] and that this number is key to characterising life on Earth and beyond [5], explaining time, matter, and more recently, even the expansion of the universe and cosmic inflation [6], according to its senior authors.
In their defence, the authors provided a single counterexample [1]. However, the proofs presented in our work hold true in the most general case [4], and our mathematical demonstration is not based on any particular implementation of $\mathbf{LZ}$ compression. Claiming that Ai is not an $\mathbf{LZ}$ scheme because its parsing differs from LZW is tantamount to claiming that LZW/LZ78 is not an $\mathbf{LZ}$ scheme because its parsings differ from LZ77, which is absurd. Additionally, there is the question of the validity of performing an experiment on a single fixed-length string without considering the rapidly converging asymptotic behaviour observed when the proper experiment is performed.
Our empirical and theoretical findings show that the assembly index is effectively an approximation of the length of the shortest statistical description of an object’s assembly process, hence a statistical upper bound of the object’s algorithmic complexity ($\mathbf{K}$).
Control experiments and comparisons of Ai to measures like LZW and other entropy-based measures—on the same datasets—and larger validation tests, were lacking, and when performed they evidence the equivalence or superiority of other traditional statistical and complexity measures compared to the assembly index, including in distinguishing biosignatures [2, 4].
Here, we further validate the fact that the assembly index converges to LZW and $\mathbf{H}$ asymptotically on the set of all strings via $\mathbf{LZ}$ compression schemes. This empirical case is just an exercise in response to the weak example the authors offered, as we have already formally proven in [4] that AT is equivalent to Shannon entropy, given that its assembly index belongs to the LZ-family encoding.
@@
Here we also prove that the Block Decomposition Method (BDM) that we introduced in the early 2010s [7], which does not shy away from combining $\mathbf{H}$ and causal local structure by way of algorithmic probability [8, 9], properly subsumes statistical compression methods such as AT and its assembly index (denoted by Ai). BDM starts by counting identical copies but also other causal operations such as reversion, complementation, inversion and other linear and non-linear transformations—known to be widely used by evolution and selection (e.g. right-left symmetry)—that AT does not consider, among other causal (and non-statistical) transformations that BDM may capture [9] but AT cannot.
We also show here that naive statistical methods are able to replicate the results reported by Assembly Theory as original, with BDM achieving a greater separation between organic and inorganic compounds when used on the same data, in particular their allegedly ‘physical’ data. This despite the fact that the ability to separate compounds is what the authors of AT provide as the main evidence in favour of its originality and relevance, claiming it as a theory of everything capable of unifying physics and biology [3]. If your claim is that their contribution is the alleged threshold, this claim has been found to be false by two other independent groups [10, 11], but we have reported it several years earlier. This does not mean that our report was false; simply that we never drew the conclusions and made the hyperbolic claims that Assembly Theory did. Unlike AT, we reported the results of our experiments as a result of applying measures of compression and algorithmic complexity to a complete database of molecular compounds [12] rather than a handful of them [3] but, more importantly, we never claimed the results to suggest that we could characterise life, matter, and time on Earth and beyond.
Although the Assembly Theory hypothesis has not been taken seriously in most academic circles and is supported mostly by amateur enthusiasts, it has been promoted by some media as a scientific breakthrough, in part driven by the senior authors’ own press releases [13, 14] and their substantial social media activities. This endless news cycle is misleading people when it comes to the progress made by complexity theory and information theory in connection to living systems. In continuing to refute hyperbolic claims made on behalf of AT, this paper represents another opportunity to correct several misconceptions about computability, algorithmic complexity, and compressibility.
2 Results
2.1 What a ZIP File Tells US About Life
In the Huffman code counterexample that the authors offer [1], they show how the assembly index (which we will denote by Ai) is suboptimal for one string, while Huffman, being equal (and smaller) on both, is not equal to Ai. We have formally demonstrated that Ai is an $\mathbf{LZ}$ compression scheme so that it converges to Shannon entropy ($\mathbf{H}$[4]. It follows then that it is equivalent to LZW and Huffman, given their convergence to Shannon entropy [15]. The authors of [1] fail to show how the assembly index does not converge to Shannon entropy in light of the mathematical proof and overwhelming empirical evidence [2]. The authors argue that being worse at compression makes Ai better, as the assembly index somehow quantifies selection while no other compression can, recycling the false argument (also partially rehashed in [16]) that Ai deals with stochastic physical, chemical or biological objects while compression can only deal with computers, deterministic systems, and bits and programs, a false and unfounded claim for which they offer no evidence.
On the one hand, their physical, chemical or biological data is represented in computable form and dealt with by a computer program implementing their Ai. On the other hand, when applying LZW– which counts for exactly the same objects they are interested in (exact repetitions and copy number)– on the data to which they have applied Ai, LZW reproduces or improves upon their own empirical results [2]. Even granting that the Ai algorithm did better at dealing with stochastic data than the quintessential theory and measure of stochasticity (Shannon entropy), it would need to be translated into a different output, preferably a better one, rather than applying a competing measure. But as we have shown, this is not the case. Moreover, as we have proven, Ai implements an $\mathbf{LZ}$ algorithm based on Shannon entropy, so if it deals with stochastic data in the same way as Shannon entropy, why would applying Ai be fundamentally different from applying Shannon entropy (as we have done and shown), albeit with extra steps.
The counterexample offered in their paper [1] amounts to the argument that Assembly Theory has the potential to implement better compression than LZW for their purposes because by design it does not provide any means to compute it efficiently. Even assuming that they had the means to better compute it because they were worse at counting copies than all existing algorithms, the assembly index would only get closer to algorithmic complexity bounded by Shannon entropy, methods that have been used for a long time, some of which take into account transformations that are relevant to selection and evolution other than concatenation and join operations (the only ones Ai can account for), such as reversion and copy variations [17]. We will discuss and demonstrate the inadequacy of such claims and counterexamples.
As recently suggested [16], we agree and suggested long ago that the discussion of AT should be discussed within the appropriate context of the theory of the theory of computation and algorithmic Information within which AT is fully subsumed [4]. But we also believe that there is no need to introduce any kind of additional stochasticity or a new type of stochastic Turing machine as suggested in the same paper, as this has been explored before [18, 19]. Furthermore, these stochastic machines are unnecessary when using a purely entropic-statistical like AT and its Ai approach already equipped to deal with stochasticity. This ability to deal with stochasticity is in no way exclusive or special to Shannon entropy, Ai or AT. Our own Block Decomposition Method (BDM) [7] deals with it by looking for all the mechanistic pathways leading to the construction of an object, as introduced over a decade ago and fully covered in the Supplementary Information.
We recognise the recent refinement of the claims made by the authors of Assembly Theory [1], who have moved from broad, unfounded assertions to the effect that AT was a generalisation of all complexity measures, including algorithmic (Kolmogorov) complexity [20], to the more modest claim that the core of Assembly Theory, its assembly index (Ai), does not instantiate an $\mathbf{LZ}$ compression algorithm that ends up simply estimating Shannon entropy. This claim is still unfounded, and furthermore, AT carries a huge freight of associated and likewise unfounded claims with no connection to selection or evolution beyond the already known connections established by way of previous results based on Shannon Entropy and compression (e.g. [12, 21, 22]).
The authors remained silent on the evidence and formal proofs provided in [2, 4] showing that their results can be easily replicated and that their assembly index calculation is equivalent to $\mathbf{H}$ by compression $\mathbf{LZ}$ (that is, it belongs to the $\mathbf{LZ}$ family of compression schemes) and not just to a specific $\mathbf{LZ}$ scheme [1].
@@ 2.1 What a ZIP File Tells US About Life
As for the efforts to provide evidence with an example indicating divergence from LZW, we show that such an example not only does not represent the general case or a correctly performed experiment, but also does not tackle any of the previous results in [2, 4], including the mathematical proofs showing equivalence with Shannon entropy via an $\mathbf{LZ}$ compression, hence with a traditional statistical compression algorithm—as also acknowledged in [16].
Another argument made by the authors in [1] relates to the suboptimality of the assembly index as a compression algorithm, which, according to the authors, the assembly index was not intended or designed to be. This constitutes two overlapping straw-man arguments.
@@
First, no requirement of compression optimality has been placed on the assembly index [4, 2]. The authors of Assembly Theory have mistakenly claimed that one of our criticisms is that their assembly index is not an optimal compression algorithm. What we actually took issue with was that it was not presented as a compression scheme, not that it was non-optimal.
In such arguments, the authors seem to misuse the concept of optimality– which is always relative to a given referential basis or context of analysis– in favour of partial counter-arguments. LZW itself is only optimal as a Shannon entropy ($\mathbf{H}$) estimator at the limit, not as an algorithmic complexity ($\mathbf{K}$) estimator, which is the index that ultimately specifies the compressiblility of an object. No compression algorithm can achieve the optimality of algorithmic complexity, so we never expected this optimality from their assembly index. Nevertheless, many types and levels of optimality can be investigated in their respective contexts, including that of counting for exact copies, which $\mathbf{LZ}$ algorithms and ultimately Shannon entropy were designed to do. Indeed, just as with the first dictionary-based compression scheme, the Huffman coding, LZW and cognates are statistical compression methods that are proven to be optimal for stochastic sources, where they approach the optimal (i.e., the minimum expected codeword length) and have as asymptotic limit that of $\mathbf{H}$.
Other examples of optimality are the approximation methods to Ai, such as the split-branch method [23] that is suboptimal with respect to what AT states as the optimal value (whose calculation is intractable) for Ai.
@@
The second part of the fallacy is that our criticism amounted to a claim that AT had the intention (or was designed) to capture the idea of optimal compression. Intentions should only play a contextual or historical role, if any. They should have no role in the scientific debate, since using arguments about intentions to evaluate the validity or soundness of a theory [1] constitutes a wishful thinking fallacy. In addition, claims regarding differences solely based on which academic field a theory is rooted in or on where the scientists’ original motivations stem from [24, 25] would also constitute a fallacy of origins.
Our criticism is that there is no empirical evidence that AT’s non-optimal compression algorithm is better than any other, or that nature or selection may favour their assembly index over any other causal coding scheme, especially in the face of our previous results replicating and improving upon their results at tasks such as separating organic from inorganic compounds [2]. The motivation and mechanism behind $\mathbf{LZ}$ algorithms is the same as that of the assembly index (Ai): to count the number of exact repetitions in an object. If Assembly Theory has pedagogical value, it is for its authors to make the case in a more restrained manner, eschewing hyperbolic claims, but it cannot be considered a fundamentally different measure than naive statistical algorithms. Assembly Theory as based on the assembly index is, therefore, not methodologically different from $\mathbf{LZ}$ algorithms in any fundamental way [4], including its possible mechanical (physical) instantiation. Nor is it a new measure, given the empirical results we have provided before in [2], and which we will further explore and extend in the present article.
It is also false that AT may explain selection or evolution. For AT to explain selection it would need to show that selection favours copy repetition in exactly the manner that Assembly Theory proposes and in no other, as such a claim would be empty if results could be achieved by any other classical naive statistical means—which, as it happens, is the case.
@@
Moreover, as we have previously discussed in [4] (later corroborated in [26]), we know that to characterise selection without invoking the relationship of an agent with its environment, such as competition for resources, is impossible for intrinsic complexity measures, such as those based on the assembly index. Furthermore, it makes no technical sense, abiotically or otherwise, as the transition to favour a certain system’s traits would still need to be explained. These conflations with selection and evolution as used in a speculative prebiotic and biological context are an unnecessary obfuscation that only misleads readers. We’ve long known that matter accumulates and clusters by the action of physical laws like gravity.
2.2 Assembly Index is a Classical Compression Algorithm
@@
The claim made in the Assembly Theory paper [1] that “the assembly index is mistaken for being what is called a ‘special case’ of algorithmic complexity without validation of the claim” shows a profound misunderstanding of the basics of information theory, both classical and algorithmic.
In contrast to mathematical fallacies that have been disseminated in some circles [16, 1] regarding information theory and theoretical computer science, no complexity measure has been found to date that is not subsumed within or is a special case of algorithmic information or Entropy. We have also formally demonstrated step-by-step not only that the assembly index and the assembly number are special-case approximations of algorithmic complexity, formalised in the language of and based on the results of algorithmic information theory (AIT) [27] (with algorithmic complexity being just one of its indexes, and usually the first index to which one is introduced), but also of a much weaker complexity measure introduced in the 1940s, Shannon entropy, and that such a special case actually belongs within the family of $\mathbf{LZ}$ algorithms.
Another argument offered is that the assembly index calculation is intractable, unlike LZW. This is true, and part of their we-are-better-because-we-are-worse argument, that is, AT’s methods, including all of those based on the assembly index, would perform better (as a complexity measure for the abiotic-to-biotic transition) than other compression algorithms because if they were a compression algorithm, then they would perform worse than other compression algorithms. In fact, the authors’ claim that their algorithm is not a compression algorithm is paradoxical, and only undermines their arguments, while bearing witness to a recurrent tendency toward obfuscation.
Not only are AT’s methods compression algorithms, but also: (i) they are subsumed within algorithmic information theory (AIT) [28, 29, 30, 27]; (ii) they are intractable according to their own proposed index calculation method, while other compression algorithms such as LZW are efficient; (iii) there is a lack of empirical evidence that they outperform other compression algorithms when used for the same ends. Given that (i) and (ii) are formally demonstrated [2, 4], a proper scientific response and debate should be focused on (iii), and not on either a straw-man argument (see Section 2.1) or a we-are-better-because-we-are-worse fallacy.
As we demonstrated in [4], such arguments do not distinguish AT’s index from those of the $\mathbf{LZ}$ family of dictionary-based compression algorithms. Regardless of whether the authors of Assembly Theory refuse to see their index as a compression scheme, suboptimal or otherwise, here we prove, again, that Assembly Theory with its assembly index is a Shannon entropy estimator of the $\mathbf{LZ}$ family, indistinguishable from popular lossless compression algorithms like ZIP.
We have shown that popular lossless compression algorithms produce the same (or better) results on the same data used by Assembly Theory, including the separation of organic from inorganic compounds [12] (albeit without the hype); and there is no evidence that Assembly Theory is closer to how selection or evolution operates than any other scheme, including other compression algorithms and complexity indexes [22].
Figure 1: Comparison of complexity values after evaluation of each index on 10,000 random permutations of the ZBC sequence with a fixed length. The Assembly Index and LZW are identical, showing convergence in permutations, with the Shannon entropy Rate aligning as string length increases. This implies that assembly index and entropy become comparable to computational compressibility measures, reflecting the consistent complexity estimates these measures provide in permutational combinatorics of molecular structures with repeated sequences. Pairwise t-test statistics between the measures were: LZW vs. Assembly Index was Not Significant (t=-0.96, p=0.37), LZW vs. BDM was significant (t=-120.95, p=2.44E-14), and assembly index (Ai) vs. BDM was significant (t=-129.91, p=1.38E-14). BDM’s greater granularity in telling apart complexity values is shown by the size of the box.

### v10

@@ Abstract
Assembly Theory (AT) and its central measure, the assembly index (Ai), represent an invaluable opportunity to address some of the most persistent and widespread conflations and misconceptions about computability and complexity theory in science. The AT defence embodies several common concurrent misconceptions that pile on each other: the belief that Turing machines impose artefactual constraints, the mischaracterisation of Kolmogorov complexity as inapplicable, and the claims around Ai as different from Shannon entropy or compression algorithms. Here we show that the new arguments advanced by the AT group in their defence, are based on misleading and incomplete experiments that, when completed, show the extent of the correlations and overlapping with popular statistical compression algorithms, conforming with the mathematical equivalence to Shannon entropy previously mathematically proved and reported, which remains undisputed. Through theoretical and empirical analysis, we show that Ai does not offer a path towards fundamental novel causal or informational insights beyond what existing statistical frameworks already offer. Rather than offering a unifying theory of life as the AT authors suggest, we argue that AT obfuscates the field and provides a cautionary example of how the accumulation of conceptual mistakes can lead to a misleading theory. Finally, we show that Ai is a particular limited case of another complexity metric based on algorithmic (Kolmogorov) complexity, consisting of decomposing an object into its causal blocks that goes beyond, and outperforms, AT.
@@
Keywords: Biosignatures, molecular complexity, Assembly Theory, algorithmic complexity, Shannon entropy, LZ compression, LZW, CTM, BDM, computability, stochasticity, invariability, data representation, Kolmogorov-Chaitin complexity, algorithmic probability, technosignatures.
1 What a ZIP File Cannot Tell About Life
Assembly Theory (AT) is a framework introduced recently that seeks to explain and quantify selection and evolution, and, consequently, life, by focussing on the minimal number of steps required to build an object of interest from its basic components. According to its authors, it emphasises historical and causal construction pathways rather than static descriptions of structure, proposing that objects with high “assembly indices” carry evidence of selection processes and cumulative history [1, 2]. By quantifying how hard it is to recreate something from simpler parts, AT claims to provide a metric for identifying signatures of life and distinguishing naturally occurring complexity from randomness. Although these descriptions sound exactly how algorithmic (Kolmogorov) complexity, Shannon entropy, and popular statistical compression approaches have tried to characterise complexity and life before, the authors insist that theirs is a different and novel approach.
@@ 1 What a ZIP File Cannot Tell About Life
According to the AT authors, the number of exact copies in an object is a measure of its selective content [2] and that this number is key to characterising life on Earth and beyond [1], explaining time, matter and more recently even the expansion of the universe and cosmic inflation, according to the authors [3].
The primary goal of this paper is to clarify some aspects of AT and their Ai under new defence arguments as published in [4]. In particular, its equivalence to compression algorithms and Shannon entropy. In doing so, we will take advantage to address common misconceptions of computability and complexity theories in science, of which AT is a quintessential example and an opportunity as a case study.
In a recent article [4], responding to some of these criticisms, putative evidence was offered to claim that AT and its Ai metric are not statistical compression algorithms or at least not popular algorithms or equivalent to Shannon entropy.
Early published findings demonstrated that Assembly Theory (AT) and a version of their Assembly Index (Ai) called molecular assembly or MA when applied to molecular complexity, performed similarly or suboptimally compared to simple statistical compression algorithms such as LZW in their own goal at separating organic from non-organic molecules, with a strong Pearson correlation of 0.874 [5]. In what follows, Ai and MA will be used interchangeably, since Ai includes MA, and MA is a specific case of Ai for molecular complexity.
In the same paper [5], it was also shown that LZW and other compression methods, including those based on Shannon entropy ($\mathbf{H}$), were able to distinguish the same organic from inorganic compounds using the AT authors’ own mass spectral data as effectively as, or better than, the assembly index. We concluded that Ai [5], and therefore AT, offered no substantial improvement over existing statistical methods for identifying biosignatures or separating living from nonliving systems, which has been a central basis of the Assembly Theory authors’ claim that their theory explained evolution and selection, and unified biology and physics [2]. However, the authors had never offered such basic control experiments in any of their previous papers, until now [4]. However, we will argue that the authors only offered a handful of examples that are misleading and incomplete, and they did not address the mathematical proof and the overwhelming evidence that their results can be all replicated by simple existing algorithms, including LZW and Shannon entropy and that there is no justification for them to believe that nature compresses objects as they propose than in any other form, but we do also have evidence that nature does not operate like AT given that AT and Ai cannot account for basic chemical and biological operations.
@@
Here we also prove that the Block Decomposition Method (BDM) that we introduced in the early 2010s [6], which does not shy away from combining $\mathbf{H}$ and causal local structure by algorithmic probability [7, 8], properly subsumes statistical compression methods such as AT and its assembly index (denoted by Ai). BDM starts by counting identical copies but also other causal operations such as reversion, complementation, inversion, and other linear and nonlinear transformations—known to be widely used by evolution and selection (e.g. right-left symmetry)—that AT does not consider, among other causal (and nonstatistical) transformations that BDM may capture [8] but AT cannot.
1.1 Preliminary Concepts
Introduced by Lempel, Ziv, and Welch in the late 1970s, the LZW algorithm that carries their initials was the first dictionary-based algorithm after Huffman’s first compression algorithm that started off the field of data compression [9]. Over time, many other algorithms, such as LZ77 and LZ78, were introduced. which in turn underlie popular compression algorithms such as ZIP, GZIP and PNG. All these algorithms are based on the same counting principle that AT and their Ai seem to suggest they have introduced. The differences between all those compression algorithms are small variations introduced for reasons such as efficiency that were not originally considered, and of which AT with their Ai is one.
A popular application of lossless compression algorithms like LZW and others in the LZ family is to estimate the algorithmic complexity (Kolmogorov or Kolmogorov-Chaitin), the accepted mathematical definition of randomness based on the length of the shortest computer programs capable of mechanistically generating an object. Formally, $\mathbf{K_{U}}(x)=\min\{|p|,U(p)=x\}$.
This metric is deeply connected to another fundamental concept in complexity theory, that of Algorithmic Probability (AP). As introduced by Solomonoff [10], AP assigns a higher likelihood of mechanistic production to objects that can be produced by shorter programs and a lower probability to objects that do not have shorter descriptions. The Coding Theorem Method (CTM) connects these two metrics using the relation $\mathbf{K}(x)\approx-\log_{2}AP(x)$ known as Coding Theorem [7] where $AP(x)$ denotes the algorithmic probability of $x$, defined as $AP(x)=\sum_{p:U(p)=x}2^{-|p|}$ where $|p|$ is the length of the programme $p$ and $U$ an optimal universal reference constructor $U$ with the length of $U$ absorbed into a constant common to all measurements [11]. The instantiation of CTM involves a (Levin) search that finds multiple (short) computer programs (or pathways) to generate $x$.
A method that extends CTM to approximate $\mathbf{K}$ via $AP$ called the Block Decomposition Method (BDM) divides larger objects into sub-blocks, estimating the complexity of each by counting the number of random programs able to reconstruct (assemble) the smaller objects, introducing a logarithmic penalty for repeated structures that effectively counts each block repetition. The result is a hybrid measure that in the ‘worse case’, approximates Shannon entropy rate in linear time by simply counting assembly blocks and compressing the object optimally according to the rules of classical information theory or, in the best case, the same method finds small generative computer programs that provide a clue on the causal mechanisms and causal content that assembles such an object, thus combining the best of both classical and causal algorithmic information theories [6].
In its standard formulation, BDM decomposes an object $X$ into sub-blocks $x_{i}$ as follows:
$\text{BDM}(X)=\sum_{i}CTM(x_{i})+\log_{2}(n_{i})$ (1)
where CTM stands for Coding Theorem Method, the search in software space of generative mechanisms explaining and generating each $x_{i}$ for each $i$ composing and re-assembling $X$ [7].
The Block Decomposition Method (BDM), introduced in 2013 [12, 13, 6], is an estimator of algorithmic information redundancies $\mathbf{K}$ based on the principle of decomposing an object into parts for which algorithms complexity estimates are aggregated. It counts exact copies in data as its most basic test for nonrandomness, but unlike purely statistical measures, BDM penalises for trivial repetitions that are also found in simple stochastic processes, such as rock formations or crystals [6]. Estimation for parts is obtained using the Coding Theorem Method (CTM) [7] based on Algorithmic Probability [10, 14] (AP) which considers all statistical and algorithmic regularities in the data.
Given that CTM can only find small programs, BDM extends its power by stitching these programs together to offer a causal explanation of the object as a whole.
In what follows, we will show that naive statistical methods are able to replicate the results reported by Assembly Theory as original. Moreover, we will show that Ai is subsumed in BDM and achieves better results across multiple domains of AT application including but not exclusive of a greater separation between organic and inorganic compounds when used on the same data as provided by the AT group, in particular what they call ‘physical’ data from their mass spectral values of molecular compounds.
2 New Defence Arguments and Results
In  [4], the authors of Assembly Theory (AT) argue that being worse at compressing objects makes Ai different, as the assembly index quantifies selection and assumes no other existing statistical compressor can, recycling the argument (also partially rehashed in [15]) that Ai can deal with stochastic physical, chemical or biological objects while compression can only deal with computers, deterministic systems, and bits and programs.
The counterexample offered in their paper [4] amounts to the argument that Assembly Theory has the potential to implement better compression than LZW or ZIP for their purposes because by design it does not provide any means to compute it efficiently.
We recognise the recent refinement of the claims made by the authors of Assembly Theory [4], who have moved from broader, unfounded assertions [16] to the more modest position that AT is a minor variation of a compression algorithm that belongs to a different (slower) time complexity class. However, unfounded claims in connection to their unique capabilities, even to characterise selection and evolution, remain and continue to be made.
2.1 AT arguments against compression based on artefacts
@@
Regarding efforts to provide evidence with an example indicating some divergence from LZW, here we will show that such an example not only does not represent the general case or a correctly performed experiment, but also does not directly address the criticisms in [5, 17], including the mathematical proof showing equivalence with Shannon entropy.
@@
Another argument made by the authors in [4] relates to the suboptimality of the assembly index as a compression algorithm. The authors of Assembly Theory have mistakenly suggested that one of our criticisms is that their assembly index is not an optimal compression algorithm. What we actually took issue with was that it was not presented as a compression scheme, not that it was nonoptimal.
In such arguments, the authors seem to misuse the concept of optimality—which is always relative to a given referential basis or context of analysis— in favour of partial counter-arguments. LZW itself is only optimal as a Shannon entropy ($\mathbf{H}$) estimator in the limit, not as an algorithmic complexity ($\mathbf{K}$) estimator.
No statistical compression algorithm can achieve the optimality of algorithmic complexity, so we never expected this optimality from their assembly index.
The second part of the misunderstanding on the part of the authors of AT is that our criticism amounted to a claim that AT had the intention (or was designed) to capture the idea of optimal compression.
@@
Intentions should not play a role in scientific debate, as using arguments about intentions to evaluate the validity or soundness of a theory or metric [4] is not a scientific discussion. Claims about differences solely based on the academic field in which a theory is rooted or on where the original motivations of scientists originate [18, 19] constitute a fallacy of origin.
One of our main criticisms is that the authors of AT have not presented any evidence that AT’s nonoptimal compression algorithm is better than any other, especially in the face of our previous results replicating and improving upon their results at tasks such as separating organic from inorganic compounds [5] on which most of their results and arguments are based on or originated from.
The mathematical motivation and mechanism behind LZ algorithms are the same as that of the assembly index (Ai), to count the number of exact repetitions in an object. If Assembly Theory has some pedagogical value, it is for its authors to make the case in a more restrained manner, eschewing the authors’ claims but it is not methodologically different from Shannon entropy and the mathematical description of LZ algorithms [17] and therefore AT with its Ai can be viewed as a method that leads to ZIP file compression that can tell little to nothing about the selection or evolution of an object beyond what Shannon entropy or compressing such an object with ZIP or GZIP tools, could not already tell.
@@
Moreover, as we have previously discussed in [17] (later corroborated in [20]), we know that characterising selection without invoking the relationship of an agent interaction as selection if a function of an agent’s environment and circumstances, such as competition for resources, is impossible for intrinsic complexity measures, such as those based on the assembly index. Furthermore, it makes no technical sense, abiotically or otherwise, as the transition to favour a certain system’s traits would still need to be explained. These conflations with selection and evolution as used in a speculative prebiotic and biological context are an unnecessary obfuscation that only misleads readers.
@@
The claim made in the Assembly Theory paper [4] that “the assembly index is mistaken for being what is called a ‘special case’ of algorithmic complexity without validation of the claim” shows a profound misunderstanding of the basics of information theory, both classical and algorithmic. While proponents of AT distinguish between the concepts of “copy number” and the “assembly index,” our analysis focuses on the latter, which is the operational core of the theory in its computational form. The notion of copy number contributes conceptually to AT but, when implemented algorithmically, results in the same behaviour as Shannon entropy algorithms and LZ-style compression schemes.
We have formally demonstrated step-by-step not only that the assembly index and the assembly number are special-case approximations of algorithmic complexity, formalised in the language of—and based on—the results of algorithmic information theory (AIT) [21].
While proponents of AT argue that Ai is intractable and thus distinct from compression algorithms like LZW, we emphasize that computational complexity (e.g., resources such as time or algorithmic cost) is separate from representational capacity. A method’s intractability does not confer additional explanatory power.
The claim that the assembly index is different from dictionary-based compression because there is no dictionary is unfounded. The very book-keeping that tracks the order and frequency of molecular substructures is the dictionary: the molecular sequence provides the alphabet, and the frequency counts instantiate the symbol—frequency table required for compression. In this formulation, the dictionary is not external but intrinsic to the data itself. This feature is not novel. Turing already showed that the boundary between programme and input disappears in universal computation, where the substrate can serve simultaneously as machine and as data. Other Turing-equivalent formalisms such as Post tag systems and the $\lambda$-calculus operate precisely in this manner, with no explicit external dictionary but with the rules and data co-encoded in the substrate. Far from distinguishing Ai from compression, this places it squarely within the standard mechanics of information representation as understood across multiple models of computation.
2.2 Empirical Replication and Correlation Analysis
2.2.1 Analysis on Synthetic Strings